xXMarcelieXx:

I forgot my algebra in this part... why did they separate the fraction.. i thought we would only separate it when the denominator had a variable by itself..

xXMarcelieXx:

1 attachment
xXMarcelieXx:

@Zepdrix

Zepdrix:

That's just how fractions work :U\[\large\rm \frac{a+b}{stuff}\quad=\quad \frac{a}{stuff}+\frac{b}{stuff}\]

Zepdrix:

\[\large\rm \frac{1+(-4v)}{stuff}\quad=\quad\frac{1}{stuff}+\frac{-4v}{stuff}\]

xXMarcelieXx:

i thought the denominator would have to some type of multiplication

Zepdrix:

I don't understand what you mean D:

Zepdrix:

\[\large\rm \frac{2+3}{7}=\frac{2}{3}+\frac{3}{7}\] Two-sevenths and three-sevenths is 5 sevenths. The bottom doesn't change when you split it into pieces. It's still 2 of 7 and 3 and 7

Zepdrix:

and 3 of 7*

Zepdrix:

Are you confusing this basic fraction math with partial fraction decomposition maybe?

xXMarcelieXx:

oh wait nvm i got it ;p

xXMarcelieXx:

yeh i was getting confused by that for ex 

Zepdrix:

Ya you gotta treat the denominator as a whole thingy-ma-jigger :x The bottom can not be broken up.